# FactReasoner: A Probabilistic Approach to Long-Form Factuality Assessment for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have demonstrated vast capabilities on generative tasks in recent years, yet they struggle with guaranteeing the factual correctness of the generated content. This makes these models unreliable in realistic situations where factually accurate responses are expected. In this paper, we propose FactReasoner, a new factuality assessor that relies on probabilistic reasoning to assess the factuality of a long-form generated response. Specifically, FactReasoner decomposes the response into atomic units, retrieves relevant contexts for them from an external knowledge source, and constructs a joint probability distribution over the atoms and contexts using probabilistic encodings of the logical relationships (entailment, contradiction) between the textual utterances corresponding to the atoms and contexts. FactReasoner then computes the posterior probability of whether atomic units in the response are supported by the retrieved contexts. Our experiments on labeled and unlabeled benchmark datasets demonstrate clearly that FactReasoner improves considerably over state-of-the-art prompt-based approaches in terms of both factual precision and recall.
@@
Radu Marinescu1, Debarun Bhattacharjya1, Junkyu Lee1, Tigran Tchrakian1, Javier Carnerero Cano1, Yufang Hou1,2, Elizabeth Daly1, Alessandra Pascale1, 1IBM Research, 2IT:U - Interdisciplinary Transformation University Austria Correspondence: radu.marinescu@ie.ibm.com
@@ 1 Introduction
Most modern approaches for assessing the factuality of LLM-generated long-form responses such as FactScore Min et al. (2023), VeriScore Song et al. (2024) and others Wei et al. (2024); Bayat et al. (2025) are prompt-based approaches and consist of three main stages: 1) the response is decomposed into a set of atomic units (facts or claims) which are subsequently revised or decontextualized to make them self-contained; 2) relevant evidence (or context) is retrieved for each atomic unit from an external knowledge source such as Wikipedia, and 3) each atomic unit is evaluated against the retrieved context to determine whether it is supported (factually correct) or not and a factuality score is calculated for the response. These approaches sometimes struggle due to conflicting information between the model’s internal knowledge and conflicting information within the retrieved contexts themselves. Therefore, they typically assume that the pieces of information retrieved do not conflict or overlap with each other Min et al. (2023).
@@
In this paper, we provide a new perspective on long-form factuality assessment that departs from the prompt-based approach, especially in the evaluation stage of the assessment. Specifically, we propose a novel factuality assessor called FactReasoner that also decomposes the response into atomic units and retrieves the relevant contexts for them from an external knowledge source. However, instead of prompting another LLM to evaluate the atoms against the retrieved evidence, FactReasoner computes the probability of each atom being supported by reasoning over a graphical model that represents a joint probability distribution over the atoms and the retrieved contexts. The graphical model is constructed using probabilistic encodings of the entailment and contradiction relationships between the natural language utterances corresponding to the atoms and contexts. Furthermore, FactReasoner makes no assumptions regarding the existence of any conflicting information within the retrieved contexts.
We conduct an extensive empirical evaluation on labeled and unlabeled benchmark datasets for long-form factuality and compare against several state-of-the-art prompt-based approaches using open-source LLMs. Our results demonstrate clearly that FactReasoner improves significantly over its competitors in terms of factual precision and recall. We show that exploiting the logical relationships between atoms and all retrieved contexts, as well as between the contexts themselves, allows FactReasoner to identify correctly considerably more supported atoms than the competing approaches.
@@ 2 Background
In this section, we provide preliminaries on probabilistic graphical models and long-form factuality assessment for large language models.
Graphical Models.
A graphical model is a tuple $\mathcal{M}=\langle\mathbf{X},\mathbf{D},\mathbf{F}\rangle$, where $\mathbf{X}=\{X_{1},\ldots,X_{n}\}$ is a set of variables, $\mathbf{D}=\{D_{1},\ldots,D_{n}\}$ is the set of their finite domains of values and $\mathbf{F}=\{f_{1},\ldots,f_{m}\}$ is a set of discrete positive real-valued functions. Each function $f_{i}$ (also called factor) is defined on a subset of variables $\mathbf{S}_{i}\subseteq\mathbf{X}$ called its scope and denoted by $vars(f_{i})$. The model $\mathcal{M}$ defines a factorized probability distribution on $\mathbf{X}$: $P(\mathbf{x})=\frac{1}{Z}\prod_{j=1}^{m}f_{j}(\mathbf{x})$ where $Z=\sum_{\mathbf{x}\in\Omega(\mathbf{X})}\prod_{j=1}^{m}f_{j}(\mathbf{x})$ is the normalization constant $Z$ is known as the partition function and $\Omega(\mathbf{X})$ denotes the Cartesian product of the variables domains Koller and Friedman (2009).
A common inference task over graphical models is to compute the posterior marginal distributions over all variables. Namely, for each variable $X_{i}\in\mathbf{X}$ and domain value $x_{i}\in D_{i}$, compute: $P(x_{i})=\sum_{\mathbf{x}\in\Omega(\mathbf{X})}\delta_{x_{i}}(\mathbf{x})\cdot P(\mathbf{x})$, where $\delta_{x_{i}}(\mathbf{x})$ is $1$ if $X_{i}$ is assigned $x_{i}$ in $\mathbf{x}$ and 00 otherwise.
Long-Form Factuality.
Let $y$ be the long-form response generated by an LLM to a query $x$. Following prior work Min et al. (2023); Song et al. (2024); Wei et al. (2024), we assume that $y$ can be decomposed into a set of $n$ atomic units (or atoms) that can be either true or false, denoted by $\mathcal{A}_{y}=\{a_{1},a_{2},\ldots a_{n}\}$. An atomic unit $a_{i}\in\mathcal{A}_{y}$ is defined as a short sentence conveying one piece of information. Furthermore, given an external knowledge source $\mathcal{C}$ [^1] For example, $\mathcal{C}$ could be Wikipedia, Google Search, or a collection of documents embedded into a vector database., we say that an atomic unit $a_{i}\in\mathcal{A}_{y}$ is supported by $\mathcal{C}$ if there exists at least one piece of information in $\mathcal{C}$ (e.g., a passage) called a context that undebatably supports $a_{i}$. Otherwise, we say that the atomic unit is not supported. The factual precision $Pr(y)$ of the response $y$ with respect to a knowledge source $\mathcal{C}$ is defined as: $Pr(y)=\frac{S(y)}{|\mathcal{A}_{y}|}$, where $S(y)=\sum_{i=1}^{n}\mathbb{I}[a_{i}\textrm{ is supported by }\mathcal{C}]$ is the number of supported atomic units. Furthermore, the notion of factual recall up to the $K$-th supported atomic unit denoted by $R_{K}(y)$ is given by: $R_{K}(y)=\min(\frac{S(y)}{K},1)$. Finally, an $F_{1}$ measure for long-form factuality denoted by $F1@K$ can be defined as: $F_{1}@K(y)=\frac{2\cdot Pr(y)\cdot R_{K}(y)}{Pr(y)+R_{K}(y)}$ if $S(y)>0$, and 00 otherwise Wei et al. (2024).
@@ 3 The FactReasoner Assessor
In this section, we present FactReasoner, a novel long-form factuality assessor that uses probabilistic reasoning to assess the factuality of the generated response with respect to an external knowledge source $\mathcal{C}$. Specifically, FactReasoner builds a graphical model that represents a joint probability distribution over the atoms of the response and their relevant contexts in $\mathcal{C}$, and subsequently computes for each atom $a_{i}$ the posterior marginal probability distribution $P(a_{i})$ representing the probability of $a_{i}$ being true (or supported) with respect to the information available in $\mathcal{C}$.
@@ 3.1 A Graphical Models Based Approach
Let $y$ be the long-form response generated by an LLM for the input query $x$, and let $\mathcal{A}_{y}=\{a_{1},\ldots,a_{n}\}$ be the set of $n$ atomic units corresponding to $y$. For simplicity, but without loss of generality, we restrict ourselves to atomic units that are either facts or claims Song et al. (2024). In addition, let $\mathcal{C}_{y}=\{c_{1},\ldots,c_{m}\}$ be a set of $m$ contexts relevant to $y$’s atoms that were retrieved from an external knowledge source $\mathcal{C}$. We make no assumptions about these contexts, namely they may be overlapping and/or contradicting each other, which is often the case in realistic situations.
@@
Variables.
We associate each atom $a_{i}\in\mathcal{A}_{y}$ and context $c_{j}\in\mathcal{C}_{y}$ with a bi-valued variable denoted by either $A_{i}$ (for atoms) or $C_{j}$ (for contexts). Therefore, we have that $\mathbf{X}=\mathbf{X}_{a}\cup\mathbf{X}_{c}$ where $\mathbf{X}_{a}=\{A_{1},\ldots,A_{n}\}$ and $\mathbf{X}_{c}=\{C_{1},\ldots,C_{m}\}$, respectively. The domains of the variables contain the values true and false indicating whether the corresponding atom or context is true or false. For simplicity, we use $a_{i}$ and $\neg a_{i}$ (resp. $c_{j}$ and $\neg c_{j}$) to denote the value assignments $A_{i}=true$ and $A_{i}=false$ (resp. $C_{j}=true$ and $C_{j}=false$).
Priors.
For each variable $A_{i}\in\mathbf{X}_{a}$ (resp. $C_{j}\in\mathbf{X}_{c}$) we consider a unary factor denoted by $f(A_{i})$ (resp. $f(C_{j})$) representing the prior belief about the truthfulness of the corresponding atom (resp. context). Since we make no assumptions about the response, we set $f(a_{i})=0.5$ and $f(\neg a_{i})=0.5$, respectively. In contrast, the external knowledge source $\mathcal{C}$ is assumed to be reliable and therefore the retrieved contexts have high probability of being true (e.g., $f(c_{j})=0.99$). Note that if a context is retrieved from a less reliable source then its prior probability can be set to a smaller value.
Relationships.
In addition, we also consider binary factors denoted by $f(A_{i},C_{j})$ and $f(C_{j},C_{k})$, defined on atom-context variable pairs as well as pairs of context variables. These factors are probabilistic representations of the logical relationships between the natural language utterances corresponding to the context and atom variables. For our purpose, we use a relation model $p_{\theta}(\cdot|t,t^{\prime})$ to predict the most likely logical relationship between an ordered pair of natural language utterances from the choices {none, entail, contradict, equivalence} [^2] The “equivalence” relationship is formed if entailment is predicted for both orderings of the utterances. The “none” relationship corresponds to neutrality meaning that the two utterances are not related to each other.. The relation model can be any pre-trained BERT or LLM Liu et al. (2019); Touvron et al. (2023).
@@
For instance, if $r^{*}$ corresponds to entailment and $(X,Y)$ is a context-atom pair then the context supports the atom. Alternatively, if $r^{*}$ is a contradiction for the same $(X,Y)$ pair then the context contradicts the atom. Finally, for BERT-based relation models, the probability $p^{*}$ is given together with the predicted relationship $r^{*}$, whereas for instructed LLM-based relation models we can obtain $p^{*}$ by applying any uncertainty quantification method Lin et al. (2024); Gao et al. (2024). We use a simple white-box method that calculates $p^{*}$ as the probability of the “entailment" or “contradiction" tokens produced by the model.
@@ 3.2 Inference and Factuality Assessment
The graphical model $\mathcal{M}=\langle\mathbf{X},\mathbf{D},\mathbf{F}\rangle$ we just defined in the previous section represents a joint probability distribution over the set of atoms and relevant externally retrieved contexts. Therefore, we can use any probabilistic inference algorithm to compute the posterior marginal distribution $P(A_{i})$ for each atom $A_{i}\in\mathcal{A}_{y}$ Pearl (1988); Koller and Friedman (2009). Specifically, in our experiments, we use an approximate variational inference algorithm called Weighted Mini-Buckets Liu and Ihler (2011) to compute the marginals.
@@
In addition to the factual precision $Pr(y)$ and $F_{1}@K$ measures, we define a new entropy-based factuality measure called $\mathcal{E}(y)$ that leverages the posterior probabilities of response $y$’s atoms:
$\mathcal{E}=\frac{1}{n}\sum_{i=1}^{n}-P(a_{i})\cdot\log P(a_{i})$ (1)
@@
Clearly, if all atoms in $\mathcal{A}_{y}$ have posterior probability $P(a_{i})=0.5$, there is virtually no external information to support or contradict the atoms (we refer to these atoms as undecided atoms) then $\mathcal{E}(y)=0.150515$. On the other hand, if all atoms are true with absolute certainty ($P(a_{i})=1$), then $\mathcal{E}(y)=0$ and if all atoms are false with absolute certainty then $\mathcal{E}(y)=\infty$. Therefore, when $\mathcal{E}(y)$ is closer to 00 the response is more truthful.
3.3 The FactReasoner Pipeline and Variants

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable success in generative tasks, yet they often fall short in ensuring the factual accuracy of their outputs this limiting their reliability in real-world applications where correctness is critical. In this paper, we present FactReasoner, a novel factuality assessment framework that employs probabilistic reasoning to evaluate the truthfulness of long-form generated responses. FactReasoner decomposes a response into atomic units, retrieves relevant contextual information from external knowledge sources, and models the logical relationships (e.g., entailment, contradiction) between these units and their contexts using probabilistic encodings. It then estimates the posterior probability that each atomic unit is supported by the retrieved evidence. Our experiments on both labeled and unlabeled benchmark datasets demonstrate that FactReasoner often outperforms state-of-the-art prompt-based methods in terms of factual precision and recall.
@@
Radu Marinescu1, Debarun Bhattacharjya1, Junkyu Lee1, Tigran Tchrakian1, Javier Carnerero Cano1, Yufang Hou1,2, Elizabeth Daly1, Alessandra Pascale1, 1IBM Research, 2IT:U - Interdisciplinary Transformation University Austria, Correspondence: radu.marinescu@ie.ibm.com
@@ 1 Introduction
Most modern approaches for assessing the factuality of LLM-generated long-form responses such as FactScore Min et al. (2023), VeriScore Song et al. (2024) and others Wei et al. (2024); Bayat et al. (2025) are prompt-based approaches and consist of three main stages: 1) the response is decomposed into a set of atomic units (facts or claims) which are subsequently revised or decontextualized to make them self-contained; 2) relevant evidence (or context) is retrieved for each atomic unit from an external knowledge source such as Wikipedia, and 3) each atomic unit is evaluated against the retrieved context to determine whether it is supported (factually correct) or not and a factuality score is calculated for the response. These approaches often struggle due to conflicting information between the model’s internal knowledge and conflicting information within the retrieved contexts themselves Min et al. (2023); Song et al. (2024).
@@
In this paper, we introduce a new perspective on long-form factuality assessment that moves beyond traditional prompt-based approaches, particularly during the evaluation phase. We propose a novel factuality assessor, FactReasoner, which decomposes a response into atomic units and retrieves relevant contextual evidence for each atom from an external knowledge source. Unlike prior methods that rely on prompting a language model to evaluate these atoms against the retrieved evidence, FactReasoner estimates the probability of each atom being supported by reasoning over a graphical model. This model encodes a joint probability distribution over the atoms and their associated contexts, constructed using probabilistic representations of entailment and contradiction relationships between the natural language utterances of the atoms and the retrieved contexts.
FactReasoner addresses three important limitations of the existing prompt-based approaches:
Context Relevance Across Atoms: In multi-atom responses, contexts retrieved for one atom can be relevant – either supportive or contradictory – to another atom. Prompt-based methods struggle with this, as they require saturating the model’s context window with all of the retrieved information. FactReasoner overcomes this limitation using a compact probabilistic representation (i.e., a graphical model) of the relationships between all atoms in the response and all of the retrieved contexts.
Handling Conflicting Contexts: Sometimes, contexts retrieved for different atoms may contradict each other. FactReasoner can leverage these contradictions effectively and in a principled manner by reasoning over their probabilistic encodings which often leads to improved performance.
Leveraging LLM Strengths in NLI Tasks: LLMs excel at natural language inference tasks such as entailment and contradiction. FactReasoner builds on this strength by framing factuality assessment as a composition of these simpler tasks.
We conduct an extensive empirical evaluation on well established labeled and unlabeled benchmark datasets for long-form factuality and compare against state-of-the-art prompt-based approaches using open-source LLMs. Our results demonstrate clearly that FactReasoner improves significantly over its competitors in terms of factual precision and recall. We show that exploiting the relationships between the atoms and retrieved contexts, as well as between the contexts themselves, allows FactReasoner to identify correctly considerably more supported atoms than the competing prompt-based approaches.
@@ 2 Background
We begin by providing background on graphical models and long-form factuality for LLMs.
Graphical Models. A graphical model is a tuple $\mathcal{M}=\langle\mathbf{X},\mathbf{D},\mathbf{F}\rangle$, where $\mathbf{X}=\{X_{1},\ldots,X_{n}\}$ is a set of variables, $\mathbf{D}=\{D_{1},\ldots,D_{n}\}$ is the set of their finite domains of values and $\mathbf{F}=\{f_{1},\ldots,f_{m}\}$ is a set of discrete positive real-valued functions. Each function $f_{i}$ (also called factor) is defined on a subset of variables $\mathbf{S}_{i}\subseteq\mathbf{X}$ called its scope and denoted by $vars(f_{i})$. The model $\mathcal{M}$ defines a factorized probability distribution on $\mathbf{X}$: $P(\mathbf{x})=\frac{1}{Z}\prod_{j=1}^{m}f_{j}(\mathbf{x})$ where $Z=\sum_{\mathbf{x}\in\Omega(\mathbf{X})}\prod_{j=1}^{m}f_{j}(\mathbf{x})$ is the normalization constant $Z$ also known as the partition function and $\Omega(\mathbf{X})$ denotes the Cartesian product of the variables domains Koller and Friedman (2009).
A common inference task over graphical models is to compute the posterior marginal distributions over all variables. Namely, for each variable $X_{i}\in\mathbf{X}$ and domain value $x_{i}\in D_{i}$, compute: $P(x_{i})=\sum_{\mathbf{x}\in\Omega(\mathbf{X})}\delta_{x_{i}}(\mathbf{x})\cdot P(\mathbf{x})$, where $\delta_{x_{i}}(\mathbf{x})$ is 111 if $X_{i}$ is assigned $x_{i}$ in $\mathbf{x}$ and 0 otherwise.
Long-Form Factuality. Let $y$ be the long-form response generated by an LLM to a query $x$. Following prior work Min et al. (2023); Song et al. (2024); Wei et al. (2024), we assume that $y$ can be decomposed into a set of $n$ atomic units (or atoms) that can be either true or false, denoted by $\mathcal{A}_{y}=\{a_{1},a_{2},\ldots a_{n}\}$. An atomic unit $a_{i}\in\mathcal{A}_{y}$ is defined as a short sentence conveying one piece of information. Furthermore, given an external knowledge source $\mathcal{C}$ [^1] For example, $\mathcal{C}$ could be Wikipedia, Google Search, or a collection of documents embedded into a vector database., we say that an atomic unit $a_{i}\in\mathcal{A}_{y}$ is supported by $\mathcal{C}$ if there exists at least one piece of information in $\mathcal{C}$ (e.g., a passage) called a context that undebatably supports $a_{i}$. Otherwise, we say that the atomic unit is not supported. The factual precision $Pr(y)$ of the response $y$ with respect to a knowledge source $\mathcal{C}$ is defined as: $Pr(y)=\frac{S(y)}{|\mathcal{A}_{y}|}$, where $S(y)=\sum_{i=1}^{n}\mathbb{I}[a_{i}\textrm{ is supported by }\mathcal{C}]$ is the number of supported atomic units. Furthermore, the notion of factual recall up to the $K$-th supported atomic unit denoted by $R_{K}(y)$ is given by: $R_{K}(y)=\min(\frac{S(y)}{K},1)$. Finally, an $F_{1}$ measure for long-form factuality denoted by $F1@K$ can be defined as: $F_{1}@K(y)=\frac{2\cdot Pr(y)\cdot R_{K}(y)}{Pr(y)+R_{K}(y)}$ if $S(y)>0$, and 0 otherwise Wei et al. (2024).
@@ 3 The FactReasoner Assessor
In this section, we present FactReasoner, a novel long-form factuality assessor that leverages probabilistic reasoning to assess the factuality of the generated response with respect to an external knowledge source $\mathcal{C}$. Specifically, FactReasoner constructs a graphical model that captures a joint probability distribution over the atomic units in the response and their corresponding contexts in $\mathcal{C}$. For each atom $a_{i}$, it then computes the posterior marginal probability distribution $P(a_{i})$, which quantifies the likelihood that $a_{i}$ is true (or supported) given the information available in $\mathcal{C}$.
@@ 3.1 A Graphical Models Based Approach
Let $y$ be the long-form response generated by an LLM for the input query $x$, and let $\mathcal{A}_{y}=\{a_{1},\ldots,a_{n}\}$ be the set of $n$ atomic units corresponding to $y$. For simplicity, but without loss of generality, we restrict ourselves to atomic units that are either facts or claims Song et al. (2024). In addition, let $\mathcal{C}_{y}=\{c_{1},\ldots,c_{m}\}$ be a set of $m$ contexts relevant to $y$’s atoms that were retrieved from an external knowledge source $\mathcal{C}$. We make no assumptions about these contexts, namely they may be overlapping and/or contradicting each other, which is often the case in realistic scenarios.
@@
Variables. We associate each atom $a_{i}\in\mathcal{A}_{y}$ and context $c_{j}\in\mathcal{C}_{y}$ with a bi-valued variable denoted by either $A_{i}$ (for atoms) or $C_{j}$ (for contexts). Therefore, we have that $\mathbf{X}=\mathbf{X}_{a}\cup\mathbf{X}_{c}$ where $\mathbf{X}_{a}=\{A_{1},\ldots,A_{n}\}$ and $\mathbf{X}_{c}=\{C_{1},\ldots,C_{m}\}$, respectively. The domains of the variables contain the values true and false indicating whether the corresponding atom or context is true or false. For simplicity, we use $a_{i}$ and $\neg a_{i}$ (resp. $c_{j}$ and $\neg c_{j}$) to denote the value assignments $A_{i}=true$ and $A_{i}=false$ (resp. $C_{j}=true$ and $C_{j}=false$).
Priors. For each variable $A_{i}\in\mathbf{X}_{a}$ (resp. $C_{j}\in\mathbf{X}_{c}$) we consider a unary factor denoted by $f(A_{i})$ (resp. $f(C_{j})$) representing the prior belief about the truthfulness of the corresponding atom (resp. context). Since we make no assumptions about the response, we set $f(a_{i})=0.5$ and $f(\neg a_{i})=0.5$, respectively. In contrast, the external knowledge source $\mathcal{C}$ is assumed to be reliable and therefore the retrieved contexts have high probability of being true (e.g., $f(c_{j})=0.99$). Note that if a context is retrieved from a less reliable source then its prior probability can be set to a smaller value.
Relationships. In addition, we also consider binary factors denoted by $f(A_{i},C_{j})$ and $f(C_{j},C_{k})$, defined on atom-context variable pairs as well as pairs of context variables. These factors are probabilistic representations of the logical relationships between the natural language utterances corresponding to the context and atom variables. For our purpose, we use a relation model $p_{\theta}(\cdot|t,t^{\prime})$ to predict the most likely logical relationship between an ordered pair of natural language utterances from the choices {none, entail, contradict, equivalence} [^2] The “equivalence” relationship is formed if entailment is predicted for both orderings of the utterances. The “none” relationship corresponds to neutrality meaning that the two utterances are not related to each other.. The relation model can be any pre-trained BERT or LLM Liu et al. (2019); Touvron et al. (2023).
@@
[…]
@@ 3.2 Inference and Factuality Assessment
The graphical model $\mathcal{M}=\langle\mathbf{X},\mathbf{D},\mathbf{F}\rangle$ we just defined in the previous section represents a joint probability distribution over the set of atoms and relevant externally retrieved contexts. Therefore, we can use any probabilistic inference algorithm to compute the posterior marginal distribution $P(A_{i})$ for each atom $A_{i}\in\mathcal{A}_{y}$ Pearl (1988); Koller and Friedman (2009). Specifically, in our experiments, we use an approximate variational inference algorithm called Weighted Mini-Buckets Liu and Ihler (2011) to compute the marginals. The algorithm is extremely efficient in practice with running times less than 0.05 seconds on all of our benchmarks.
@@
In addition to the factual precision $Pr(y)$ and $F_{1}@K$ measures, we define a new entropy inspired factuality measure called $\mathcal{E}(y)$ that leverages the posterior probabilities of response $y$’s atoms:
$\mathcal{E}=\frac{1}{n}\sum_{i=1}^{n}-P(a_{i})\cdot\log P(a_{i})\vskip-1.99997pt$ (1)
@@
Clearly, if all atoms in $\mathcal{A}_{y}$ have posterior probability $P(a_{i})=0.5$, there is virtually no external information to support or contradict the atoms (we refer to these atoms as undecided atoms) then $\mathcal{E}(y)=0.150515$. On the other hand, if all atoms are true with absolute certainty ($P(a_{i})=1$), then $\mathcal{E}(y)=0$ and if all atoms are false with absolute certainty then $\mathcal{E}(y)=\infty$. Therefore, when $\mathcal{E}(y)$ is closer to 0 the response is more truthful.
3.3 The FactReasoner Pipeline